OpenStudy (anonymous):

find sin1/2B, when cosB = 12/13

OpenStudy (mertsj):

\[\cos 2B=2\cos ^2B-1=2(\frac{12}{13})^2-1=\frac{119}{169}\]

OpenStudy (mertsj):

\[\cos ^22B+\sin ^22B=1=\frac{119}{169}^2+\sin ^22B=1\]

OpenStudy (mertsj):

\[\sin ^22B=1-\frac{14161}{28561}=\frac{14400}{28561}\]

OpenStudy (mertsj):

\[\sin 2B=\frac{120}{169}\]

OpenStudy (mertsj):

\[\frac{1}{\sin 2B}=\frac{169}{120}\]

OpenStudy (anonymous):

B=1/(cos(12/13)) =0.39479 therefore, (sin1)/2(1/(cos(12/13))) =1.0657